What must be the price of a $2,000 bond with a 6.2​% coupon​ rate, annual​ coupons, and 30 years to maturity if YTM is 8.5% ​APR?

A.$ 1,505.65

B. $1,204.52

C. $2,107.90

D. $1,806.77
Answer #1

Coupon = (0.062 * 1000) = 124

Price = Coupon * [1 - 1 / (1 + r)n] / r + FV / (1 + r)n

Price = 124 * [1 - 1 / (1 + 0.085)30] / 0.085 + 2000 / (1 + 0.085)30

Price = 124 * [1 - 0.086518] / 0.085 + 173.036551

Price = 124 * 10.746847 + 173.036551

Price = $1,505.65